[clas_members] [Revised Logentry] RC Daily summary - Monday

daria at jlab.org daria at jlab.org
Mon Sep 12 13:20:02 EDT 2022

Logentry Text:
Dear RG-C enthusiasts,

It's all beginning to unravel a little. 

Yesterday's update took us to the target anneal (30 min in liquid argon), followed by polarisation with the microwaves at the newly-found "optimal" frequency. Finding it is a finicky process and the frequency had to be adjusted several times as the polarisation (and later dose) increased. When NMR was reading -12% (at 3.40 pm), we sent 7 nA on the target and concurrently started checking and adjusting the raster. 

There was an, ahem, glitch with the raster controls when I accidentally put an embarrassingly large number in the x-offset window -- the x-raster shut down but was brought back immediately by Eugene. 

Optimised raster, although it looks like it could still use a nudge up in y, we will try this at the start of the next run: https://logbooks.jlab.org/entry/4040334

Continued taking data and watching and tweaking the microwave frequency during the evening and the night (thank you, Victoria, for the middle-of-night target rescues!). 
At 4pm it got to -13.4%. At 5.20 pm, -19%. Run 16945 was the process of polarisation increasing and raster optimisation.

Just as things seemed to be going well for us, MCC started having trouble. Lost beam for ~30min for their optimisation around 7.45pm, then again at midnight for 5h30 when a breaker died in the accelerator. The polarisation levelled out as it was no longer being irradiated: https://logbooks.jlab.org/entry/4040461

Since the beam returned just before dawn, we have been taking production data with negative target polarisation, half-wave plate still OUT. Polarisation increasing, currently above 29% (negative).

Mohammad keeps cooking and Harut, Raffaella and Noemie have been studying the polarisation off-line in the meantime. The method of forming an asymmetry on-line from helicity-latched triggers looked hopeful but is still inconclusive. 

My (and Larry's) RC sreports at today's Hall B meeting: https://clasweb.jlab.org/group_meeting_minutes/Physics/talks_2022/rc-report-09122022.pdf

Let's have the RC meeting at 2pm today as usual!

Your (today) slightly frazzled RC.

******** Stats since the start of ND3 with ELMO ********
Half-wave plate OUT, target polarisation POSITIVE:

16902 - 16908: details in an earlier RC summary.

TOTAL triggers from "good" runs: 467 M (peak pol. about 37% according to NMR).

Half-wave plate IN, target polarisation POSITIVE:

16910 - 16922: details in an earlier RC summary.

TOTAL triggers from "good" runs: 498 M (slightly lower target pol than the set with OUT half-wave plate).

Half-wave plate IN, target polarisation NEGATIVE:

16923 - 16832: details in the previous RC summary.

TOTAL triggers from "good" runs: 638 M. (peak NMR pol. about -31%)

Half-wave plate OUT, target polarisation NEGATIVE:

16933 - 16944: details in the previous RC summary.

TOTAL triggers from "good" runs: 335 M. (pol. dropping during this stage).

Target ANNEALED, Half-wave plate still OUT, target polarisation again NEGATIVE:

16945: 78.4 M (raster optimisation, pol. increasing from 12% to 19% -- not for production).
16946: 26.6 M
16947: 67.3 M (DAQ crashed at end). 
16948: 17.8 M
16951: 100.2 M (NMR reading about -28% at the end, polarisation still increasing).
16952: running... currently 94M.

TOTAL triggers from "good" runs so far, with NMR pol. reading > 20% (and steadily increasing): 306 M. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4040550
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220912/ad1febeb/attachment-0001.html>

More information about the clas_members mailing list